LASER INDUCED FLUORESCENCE SPECTROSCOPY OF BORON CARBIDE

\underline{A. S-C. CHEUNG}, Y.W. NG, AND H.F. PANG , Department of Chemistry, The University of Hong Kong, Pokfulam Road, Hong Kong..

Laser induced fluorescence spectrum of boron carbide (BC) between 490 and 560 nm has been recorded and analyzed. Gas-phase BC molecule was produced by the reaction of B_{2}H_{6} and CH_{4} in the presence of magnesium atom from laser ablation process. The (0, 0), (1, 0), and (2, 0) bands of the B^{4}Σ^{−} - X^{4}Σ^{−} transition were recorded and rotationally analyzed. Spectra of both isotopes: ^{10}BC and ^{11}BC were observed. Equilibrium molecular constants for the B^{4}Σ^{−} and the X^{4}Σ^{−} states for both isotopes were determined. A comparison of the determined gas-phase molecular constants with those obtained using matrix isolation spectroscopy and the theoretical calculations will be presented. Financial support from the Research Grants Council of the Hong Kong Special Administrative Region, China (Project No. HKU 701008P) is gratefully acknowledged.